# BEST-RQ-Based Self-Supervised Learning for Whisper domain adaptation

### v1

@@ Abstract
Automatic Speech Recognition (ASR) systems, despite large multilingual training, struggle in out-of-domain and low-resource scenarios where labeled data is scarce. We propose BEARD (BEST-RQ Encoder Adaptation with Re-training and Distillation), a novel framework designed to adapt Whisper’s encoder using unlabeled data. Unlike traditional self-supervised learning methods, BEARD uniquely combines a BEST-RQ objective with knowledge distillation from a frozen teacher encoder, ensuring the encoder’s complementarity with the pre-trained decoder. Our experiments focus on the ATCO2 corpus from the challenging Air Traffic Control (ATC) communications domain, characterized by non-native speech, noise, and specialized phraseology. Using about 5,000 hours of untranscribed speech for BEARD and 2 hours of transcribed speech for fine-tuning, the proposed approach significantly outperforms previous baseline and fine-tuned model, achieving a relative improvement of 12% compared to the fine-tuned model. To the best of our knowledge, this is the first work to use a self-supervised learning objective for domain adaptation of Whisper.
@@ 1 Introduction
A way to improve a speech recognition system to a new domain with little transcribed speech and a larger amount of untranscribed audio, is self-training [2]. A teacher model is first trained using the transcribed audio, and then used to create pseudo-labels for the untranscribed speech. A student model is then trained using the pseudo-labels in a supervised way. Though this method has shown to be successful [3], pseudo-labeling large amounts of data is computationally expensive and sensitive to the teacher’s model accuracy.
Self-supervised learning (SSL) methods were introduced to perform model training in the absence of annotations. SSL methods enable leveraging large amounts of unlabeled data to learn informative representations through representation learning. A decoder can then be added on top of the SSL encoder and trained in a supervised manner for ASR. wav2vec 2.0 [4] proposes using contrastive learning to predict representations of masked parts. This method, while proven effective across diverse domains [5, 6], is computationally expensive. Instead of using contrastive learning, HuBERT [7] employs k-means clustering to learn a quantizer that maps speech signals into discrete labels. It then performs BERT-style [8] prediction pre-training, by masking regions of the input. The prediction loss is applied over the masked regions. Both wav2vec 2.0 and HuBERT have been utilized to improve speech recognition for low-resource domains [9]. More recently, BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) [10] was introduced as a lighter SSL alternative for training speech encoders. The approach learns a model to predict the masked speech signals, in the form of discrete labels generated with a random-projection quantizer. The quantizer projects speech inputs with a randomly initialized matrix, and does a nearest-neighbor lookup in a randomly initialized codebook. As the quantizer is kept frozen and independent from the ASR model, this strategy remains adaptable and suitable for integration with any speech recognition framework. BEST-RQ has showed speech recognition improvements in different domains [11, 12].
@@
In this paper, we propose BEARD (BEST-RQ Encoder Adaptation with Re-training and Distillation), a framework that adapts Whisper’s encoder with unlabeled speech to improve ASR performance on new domains. Unlike traditional SSL methods, which are designed to pre-train speech encoders from scratch, BEARD makes use of Whisper’s already powerful encoder. By combining self-supervised learning and distillation, BEARD re-trains Whisper’s encoder on untranscribed speech while preserving its complementarity with the decoder. We show that BEARD gives a stronger base to fine-tune on low-resource labeled data from the same domain. Using BEARD decreases significantly the word error rate on the ATCO2 corpus compared to previous works [18, 19] and a model fine-tuning using only labeled data. Compared to previous studies, we apply an SSL objective to a middle-layer of the speech encoder and use distillation at the same time. To the best of our knowledge, this is the first work that leverages self-supervised learning to improve Whisper.
@@ 2 Proposed methodology
Table 1: WER (%) obtained with different fine-tuning methods on ATCO2. The Adaptation method column indicates the method used to adapt the ASR model. The ATCO2 for FT column indicates the quantity of ATCO2 audio used for fine-tuning. Bold numbers indicate the best result and those results which are statistically equivalent to it. * indicates results reported in the literature.
Adaptation method # trained params. ATCO2 for FT Layer $\ell$ Distillation weight $\bm{\lambda}$ WER (%)
@@
Whisper-small, BEARD + FT 244M 2 h 24 min 8 0.5 18.40
1.0 18.27
7 0.5 17.87
1.0 19.68
6 0.5 17.17
1.0 18.05
5 0.5 17.72
1.0 18.06
4 0.5 18.01
1.0 17.58
@@ 2.2 BEST-RQ
BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) [10] is a self-supervised learning approach which converts input speech signal into discrete labels. This is done with a random-projection quantizer, which projects the input signal and finds the nearest vector in a codebook. The index of that vector becomes the discrete label. During pre-training, portions of the input log-mel spectrogram are masked and fed into a speech encoder. The speech encoder is trained to predict the discrete labels of the masked parts. The random-projection quantizer is kept frozen throughout the pre-training stage. After pre-training the speech encoder, a decoder can be added on top and trained for the ASR task. BEST-RQ had first been proposed as a pre-training approach for Conformer architectures [21]. It was demonstrated that BEST-RQ can also be applied to convolution-free Transformer architectures [22].
@@ 2.3 BEST-RQ Encoder Adaptation with Re-training and Distillation (BEARD)
Using only $\mathcal{L}_{q}^{\ell}$ would lead to changes in the model up to the $\ell$-th layer, leaving the following layers untouched. In order to adapt all layers, and, at the same time, preserve the complementarity with the decoder, we introduce two distillation losses using the frozen teacher encoder (T). $\mathcal{L}_{d}^{n}$ is computed between the last layer outputs of S and T. By making sure S’s output space remains close to T’s, we assure the complementarity with the decoder is preserved. $\mathcal{L}_{d}^{\ell}$ is computed between the outputs of the $\ell$-th layer of S and the $\ell$-th layer of T, which helps the distillation process. Both $\mathcal{L}_{d}^{n}$ and $\mathcal{L}_{d}^{\ell}$ use cosine similarity, which should be maximized. We use cosine similarity over L1 or mean squared error, because it is less constraining, allowing for the encoder to adapt to the new domain. Furthermore, as using SSL requires masking parts of S’s input, the distillation losses are computed exclusively on the unmasked regions.
@@
After re-training S using BEARD, we propose adding back Whisper’s decoder on top of S. We then fine-tune them both, without BEARD, using labeled speech from the same domain as the untranscribed speech.
@@ 3.1 Dataset
We conducted our experiments using the ATCO2 dataset [^1] ATCO2 project data, ELRA catalog (http://catalog.elra.info), ISLRN : 589-403-577-685-7, ELRA ID : ELRA-S0484 [13]. It contains air traffic control communications between pilots and air traffic controllers from various airports. The speech is non-native, with high speech rate and noisy, with signal-to-noise ratios (SNR) varying from -10dB to 40dB, estimated using WADA-SNR [23]. The corpus consists of 2 parts: 5,381 hours of audio without transcription, and 4 hours of audio with human transcriptions [^2] Only one hour of transcribed audio is freely available.. We do not have any information about the number of speakers and their native language. For the self-supervised re-training stage, we employ the untranscribed part of the dataset. For the fine-tuning stage, we use the transcribed speech part. We conduct 4-fold cross-validation. For a given fold, 2 h 24 min (25,000 words) are used for the fine-tuning, 36 min (5,300 words) for validation, and 1 h (10,000 words) for testing. All the audio files of ATCO2 are sampled at 16kHz, which matches Whisper’s input requirements.
3.2 General parameters
@@
Our experiments were conducted with the Whisper-small model [20] comprising 244M parameters. Its compact size makes it suitable for deployment on a wide range of devices, including embedded systems, while remaining a high-performance ASR model. Its parameter scale (244M) is comparable to the XLS-R model [5] (300M) employed as the baseline for ATCO2 [18]. Moreover, both Whisper-small and XLS-R are end-to-end ASR systems trained on a wide variety of languages. For BEARD re-training stage, models are re-trained for a single epoch over all untranscribed data available, with a batch size of 32. We chose to run only one epoch since over 5,000 hours is a significant amount of data. We set the learning rate to 1e-5 for Whisper’s encoder and 5e-4 for the projection layer. We use a masking span of 4 frames [24]. We reduce the masking probability from 0.15 to 0.10 to ensure that a sufficient number of unmasked frames remain for the distillation. For the random-projection quantizer, we use a codebook size of 2048, which obtained the best results in our preliminary experiments. We set $\beta$ to 0.1 to down-weight $\mathcal{L}_{d}^{n}$. Higher values led to worse results. On 8 NVIDIA V100 GPUs, one epoch of BEARD re-training takes around 7 hours on the 5,381 hours of untranscribed speech. This is much shorter than creating pseudo-labels. For the fine-tuning stage, models are trained until convergence, with a batch size of 16. The learning rate is set to 1e-5. For decoding, we use greedy search for computational reasons. Our code is publicly available  [^3] https://gitlab.inria.fr/rbagat/beard.
@@
The results are reported in terms of the word error rate (WER). Early stopping is made using the WER on the validation set. The statistical significance of the results has been validated using the matched pair sentence segment test with SCTK [25] ($p=0.001$).
@@ 4.1 Baselines
The reported four baselines are shown in the first four rows of Table 1. Among these, No FT performs worst, with a WER of 63.32%. XLS-R FT reaches 19.80%, outperforming Whisper-small FT (22.79%). FT (2 h 24 min), using the same base model as Whisper-small FT, but fine-tuned with more data, achieves a WER of 19.54%, which is comparable to XLS-R FT.
@@ 4.2 BEARD results
For every experiment, we take the pre-trained Whisper-small model and we re-train its 12-layer encoder using BEARD. Then, we fine-tune both encoder and decoder on transcribed speech in a cross-validation setup. We evaluated BEARD under multiple conditions with different parameter settings: encoder layer $\ell\in\{4,5,6,7,8\}$ and weighting factor $\lambda\in\{0.5,1\}$. These values were chosen to evaluate the sensibility of the method to the choice of the layer $\ell$, and to see whether the distillation should be more balanced or not.
@@
Effect of $\bm{\lambda}$ — To assess the influence of the distillation weight $\lambda$, we evaluated two values, $\lambda\in\{0.5,1.0\}$, for each of the encoder layer $\ell$ to which BEARD was applied. As shown in Table 1, lower value of $\lambda$ leads to better performance across most layers. For layers $\ell=5$, $6$, and $7$, using $\lambda=0.5$ resulted in significantly lower WERs compared to $\lambda=1.0$. In all cases, except when BEARD is applied at layer 7 with $\lambda=1.0$, the results significantly outperform FT’s results.
@@ 5 Conclusion
In this work, we investigated whether self-supervised learning can help Whisper adapt to a new domain. We introduced BEARD, a framework that combines self-supervised learning and distillation to adapt Whisper’s encoder using unlabeled speech. The modified encoder is then fine-tuned with the decoder using a limited amount of labeled data. On the ATCO2 corpus, the best BEARD configuration significantly reduced the word error rate, with a relative improvement of 12% over fine-tuning with transcribed speech only. BEARD showed WER improvements across all SNR ranges. The ablation study demonstrated that distillation is important to keep the encoder-decoder complementarity. These results confirm that large-scale unlabeled data can be effectively exploited through self-supervision to adapt Whisper ASR model. The proposed approach can be used to adapt any encoder-decoder models to new domains. Future work will study alternative SSL objectives, use different quantity of unlabeled data (not currently tested due to computational limitations), filter the unlabeled data, and adapt larger Whisper variants.

### v2

@@ Abstract
Automatic Speech Recognition (ASR) systems, despite large multilingual training, struggle in low-resource scenarios where labeled data is scarce. We propose BEARD (BEST-RQ Encoder Adaptation with Re-training and Distillation), a novel framework designed to adapt Whisper’s encoder with unlabeled data. Unlike traditional self-supervised learning methods, BEARD uniquely combines a BEST-RQ objective with knowledge distillation from a frozen teacher encoder, ensuring the encoder’s complementarity with the pre-trained decoder. Our experiments focus on the ATCO2 corpus from the challenging Air Traffic Control (ATC) communications domain, characterized by non-native speech, noise, and specialized phraseology. Using about 5,000 hours of untranscribed speech for BEARD and 2 hours of transcribed speech for fine-tuning, the proposed approach significantly outperforms previous baseline and fine-tuned model, achieving a relative improvement of 12% compared to the fine-tuned model. To the best of our knowledge, this is the first work to use a self-supervised learning objective for domain adaptation of Whisper.
@@ 1 Introduction
A way to improve a speech recognition system to a new domain with little transcribed speech and a larger amount of untranscribed audio, is self-training [23]. A teacher model is first trained using the transcribed audio, and then used to create pseudo-labels for the untranscribed speech. A student model is then trained using the pseudo-labels in a supervised way. Though this method has shown to be successful [14], pseudo-labeling large amounts of data is computationally expensive and sensitive to the teacher’s accuracy.
Self-supervised learning (SSL) methods were introduced to perform model training in the absence of annotations. SSL methods enable leveraging large amounts of unlabeled data to learn informative representations through representation learning. A decoder can then be added on top of the SSL encoder and trained in a supervised manner for ASR. wav2vec 2.0 [3] proposes using contrastive learning to predict representations of masked parts. This method, while proven effective across diverse domains [1, 11], is computationally expensive. Instead of using contrastive learning, HuBERT [9] employs k-means clustering to learn a quantizer that maps speech signals into discrete labels. It then performs BERT-style [5] prediction pre-training, by masking regions of the input. The prediction loss is applied over the masked regions. Both wav2vec 2.0 and HuBERT have been utilized to improve speech recognition for low-resource domains [25]. The use of semi-supervised with self-supervised training has been explored for unseen domain adaptation [13, 12]. More recently, BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) [4] was introduced as a lighter SSL alternative for training speech encoders. The approach learns a model to predict the masked speech signals, in the form of discrete labels generated with a random-projection quantizer. The quantizer projects speech inputs with a randomly initialized matrix, and does a nearest-neighbor lookup in a randomly initialized codebook. As the quantizer is kept frozen and independent from the ASR model, this strategy remains adaptable and suitable for integration with any speech recognition framework. BEST-RQ has showed speech recognition improvements in different domains [24, 10].
@@
In this paper, we propose BEARD (BEST-RQ Encoder Adaptation with Re-training and Distillation), a framework that adapts Whisper’s encoder with unlabeled speech to improve ASR performance on new domains. Unlike traditional SSL methods, which are designed to pre-train speech encoders from scratch, BEARD makes use of Whisper’s already powerful encoder. By combining self-supervised learning and distillation, BEARD re-trains Whisper’s encoder on untranscribed speech while preserving its complementarity with the decoder. We show that BEARD gives a stronger base to fine-tune on low-resource labeled data from the same domain. Using BEARD decreases significantly the word error rate (WER) on the ATCO2 corpus w.r.t. previous works [26, 20] and a model fine-tuning using only labeled data. W.r.t. previous studies, we apply an SSL objective to a middle-layer of the speech encoder and use distillation at the same time. To the best of our knowledge, this is the first work that leverages self-supervised learning to improve Whisper.
@@ 2 Proposed methodology
Table 1: WER (%) obtained with different fine-tuning methods on ATCO2. The Adaptation method column indicates the method used to adapt the ASR model. The ATCO2 for FT column indicates the quantity of ATCO2 audio used for fine-tuning. For BEARD, we indicate how many hours of unlabeled audio were used in parentheses. Bold numbers indicate the best result and those results which are statistically equivalent to it according to SCTK’s matched pair sentence segment test. * indicates results reported in the literature.
Adaptation method # FT params. ATCO2 for FT Layer $\ell$ Distillation weight $\bm{\lambda}$ WER (%)
@@
Whisper-small, BEARD (5381 h) + FT 244M 2 h 24 min 8 0.5 / 1.0 18.40 / 18.27
7 0.5 / 1.0 17.87 / 19.68
6 0.5 / 1.0 17.17 / 18.05
5 0.5 / 1.0 17.72 / 18.06
4 0.5 / 1.0 18.01 / 17.58
Whisper-small, BEARD (2000 h/1000 h/500 h) + FT 244M 2 h 24 min 6 0.5 18.40 / 18.06 / 17.53
@@ 2.2 BEST-RQ
BERT-based Speech pre-Training with Random-projection Quantizer (BEST-RQ) [4] is a self-supervised learning approach which converts input speech signal into discrete labels. This is done with a random-projection quantizer, which projects the input signal and finds the nearest vector in a codebook. The index of that vector becomes the discrete label. During pre-training, portions of the input log-mel spectrogram are masked and fed into a speech encoder. The speech encoder is trained to predict the discrete labels of the masked parts. The random-projection quantizer is kept frozen throughout the pre-training stage, which makes it easier to train than wav2vec2.0 or HuBERT’s. After pre-training the speech encoder, a decoder can be added on top and trained for the ASR task. BEST-RQ had first been proposed as a pre-training approach for Conformer architectures [7]. It was demonstrated that BEST-RQ can also be applied to convolution-free Transformer architectures [8].
@@ 2.3 BEST-RQ Encoder Adaptation with Re-training and Distillation (BEARD)
Using only $\mathcal{L}_{q}^{\ell}$ would lead to changes in the model up to the $\ell$-th layer, leaving the following layers untouched. In order to adapt all layers, and, at the same time, preserve the complementarity with the decoder, we introduce two distillation losses using the frozen teacher encoder (T). $\mathcal{L}_{d}^{n}$ is computed between the last layer outputs of S and T. By making sure S’s output space remains close to T’s, we assure the complementarity with the decoder is preserved. $\mathcal{L}_{d}^{\ell}$ is computed between the outputs of the $\ell$-th layer of S and the $\ell$-th layer of T, which helps the distillation process. Both $\mathcal{L}_{d}^{n}$ and $\mathcal{L}_{d}^{\ell}$ use cosine similarity, which should be maximized. We use cosine similarity over L1 or mean squared error, because it is less constraining as it does not penalize vector norms, allowing for the encoder to adapt to the new domain. Furthermore, as using SSL requires masking parts of S’s input, the distillation losses are computed exclusively on the unmasked regions.
@@
After re-training S using BEARD, we propose adding back Whisper’s decoder on top of S. We then fine-tune them both, without BEARD, using labeled speech from the same domain as the untranscribed speech. Compared to previous works, we adapt a pre-trained supervised model. [13] proposes joint SSL/supervised training, while [12] explores SSL+FT sequentially. In our approach, we use SSL to adapt a pre-trained supervised model along with distillation.
@@ 3 Experimental settings
Dataset — We conducted our experiments using the ATCO2 dataset [^1] ATCO2 project data, ELRA catalog (http://catalog.elra.info), ISLRN : 589-403-577-685-7, ELRA ID : ELRA-S0484 [27]. It contains air traffic control communications between pilots and air traffic controllers from various airports. The speech is non-native, with high speech rate and noisy, with signal-to-noise ratios (SNR) varying from -10 dB to 40 dB, estimated using WADA-SNR [15]. The corpus consists of 2 parts: 5,381 hours of audio without transcription, and 4 hours of audio with human transcriptions [^2] Only one hour of transcribed audio is freely available.. We do not have any information about the number of speakers and their native language. For the self-supervised re-training stage, we employ the untranscribed part of the dataset. For the fine-tuning stage, we use the transcribed speech part. We conduct 4-fold cross-validation. For a given fold, 2 h 24 min (25,000 words) are used for the fine-tuning, 36 min (5,300 words) for validation, and 1 h (10,000 words) for testing. All the audio files of ATCO2 are sampled at 16 kHz, which matches Whisper’s input requirements.
@@
General parameters — Our experiments were conducted with the Whisper-small model [19] comprising 244M parameters. Its compact size makes it suitable for deployment on a wide range of devices, including embedded systems, while remaining a high-performance ASR model. Its parameter scale (244M) is comparable to the XLS-R model [1] (300M) employed as the baseline for ATCO2 [26]. Moreover, both Whisper-small and XLS-R are end-to-end ASR systems trained on a wide variety of languages. For BEARD re-training stage, models are re-trained for a single epoch over all untranscribed data available, with a batch size of 32. We chose to run only one epoch since over 5,000 hours is a significant amount of data. We set the learning rate to 1e-5 for Whisper’s encoder and 5e-4 for the projection layer. We use a masking span of 4 frames [21]. We reduce the masking probability from 0.15 to 0.10 to ensure that a sufficient number of unmasked frames remain for the distillation. For the random-projection quantizer, we use a codebook size of 2048, which obtained the best results in our preliminary experiments. We set $\beta$ to 0.1 to down-weight $\mathcal{L}_{d}^{n}$. Higher values led to worse results. On 8 NVIDIA V100 GPUs, one epoch of BEARD re-training takes around 7 hours on the 5,381 hours of untranscribed speech. This is much shorter than creating pseudo-labels. For the fine-tuning stage, models are trained until convergence, with a batch size of 16. The learning rate is set to 1e-5. For decoding, we use greedy search for computational reasons. Our code is publicly available  [^3] https://gitlab.inria.fr/rbagat/beard.
@@
The results are reported in terms of the WER. Early stopping is made using the WER on the validation set. The statistical significance of the results has been validated using the matched pair sentence segment test with SCTK [17] ($p=0.001$).
@@ 4.1 Baselines
The reported four baselines are shown in the first four rows of Table 1. Among these, No FT performs worst, with a WER of 63.32%. XLS-R FT reaches 19.80%, outperforming Whisper-small FT (22.79%). FT (2 h 24 min), using the same base model as Whisper-small FT, but fine-tuned with more data, achieves a WER of 19.54%, which is comparable to XLS-R FT. As a side experiment, we fine-tuned Whisper-small using pseudo-labels [14] from 3,500 hours of untranscribed ATCO2 audio and obtained 25.97% WER which is significantly worse than Whisper-small FT (22.79%).
@@ 4.2 BEARD results
For every experiment, we take the pre-trained Whisper-small model and we re-train its 12-layer encoder using BEARD. Then, we fine-tune both encoder and decoder on transcribed speech in a cross-validation setup. We evaluated BEARD, by using 5,381 h of unlabeled audio, under multiple conditions with different parameter settings: encoder layer $\ell\in\{4,5,6,7,8\}$ and weighting factor $\lambda\in\{0.5,1.0\}$. These values were chosen to evaluate the sensibility of the method to the choice of the layer $\ell$, and to see whether the distillation should be more balanced or not.
@@
Effect of $\bm{\lambda}$ — To assess the influence of the distillation weight $\lambda$, we explored different values. We only report $\lambda=0.5$, which is the best value we found, and $\lambda=1.0$ for comparison, for each of the encoder layer $\ell$ to which BEARD was applied. As shown in Table 1, lower value of $\lambda$ leads to better performance across most layers. For layers $\ell=5$, $6$, and $7$, using $\lambda=0.5$ resulted in significantly lower WERs compared to $\lambda=1.0$. In all cases, except when BEARD is applied at layer 7 with $\lambda=1.0$, the results largely outperform FT.
@@
Effect of data quantity for BEARD — We also applied BEARD using different amounts of unlabeled data in our best configuration, as reported in the last row of Table 1, and obtained WERs of 18.40%, 18.06% and 17.53%. This shows that BEARD is also beneficial when less unlabeled data are available.
@@ 5 Conclusion
We investigated whether self-supervised learning can help Whisper adapt to a new domain. We introduced BEARD, a framework that combines self-supervised learning and distillation to adapt Whisper’s encoder using unlabeled speech. The modified encoder is then fine-tuned with the decoder using a limited amount of labeled data. On the ATCO2 corpus, the best BEARD configuration achieved a relative 12% WER improvement over fine-tuning with transcribed speech only. BEARD showed WER improvements across all SNRs. The ablation study demonstrated that distillation is important to keep the encoder-decoder complementarity. These results confirm that large-scale unlabeled data can be effectively exploited through self-supervision to adapt Whisper ASR model. The proposed approach can be used to adapt any encoder-decoder models to new domains. Future work will study other SSL objectives, multi-layer SSL adaptation, filter the unlabeled data, and adapt a larger Whisper.